### v1

# Magnetic field generation in mergers of massive main sequence stars

Magnetic field generation in mergers of massive main sequence stars
Magnetic fields are found in many astrophysical objects, ranging from galaxy clusters to the interstellar medium of galaxies and magnetars. Strong magnetic fields are also observed in massive stars, but it is still unclear how they are generated. There are different theories for their origin: the magnetic fields could be fossil fields from star formation, they could be generated in dynamo processes during stellar evolution, or they could be generated during a merger of two stars. Here, we show how magnetic fields are generated in simulations of mergers of massive main-sequence stars, conducted with the 3D magnetohydrodynamics code Arepo. In these simulations, a and a main sequence star merge, whereby the more massive star transfers mass to the less massive one until it is finally disrupted, forming a core–disk structure. During the initial mass transfer, magnetic fields start to be amplified in the accretion stream, saturating after the disruption at an equipartition level. Correspondingly, the magnetic energy starts being present at small scales, growing to larger scales during the merger. During the merger, the magnetic pressure is much smaller than the gas pressure, thus the dynamical impact of the magnetic fields on the merger is small. After the merger, both are similar in some regions and the magnetic field shows a large-scale structure that is mostly toroidal and that is expected to be stable. Thus, our simulations show that magnetic fields in massive stars can originate from a merger of two main sequence stars.
Key Words.:
hydrodynamics – methods: numerical – stars: general1 Introduction
The merger of two stars is a common phenomenon in stellar astrophysics and may invoke compact objects such as neutron stars and black holes, but also stars during their main nuclear-burning cycles. While the former are observed, e.g., in gravitational-wave events (Abbott et al. 2016, 2017), the latter lead to outbursts that are called luminous red novae with V1309 Sco and V838 Mon being prototypical events (Munari et al. 2002; Tylenda et al. 2011). Merged stars may be rejuvenated (e.g. Hellings 1983; Braun & Langer 1995; Dray & Tout 2007; Schneider et al. 2016) and can appear as blue stragglers in star clusters and stellar associations (e.g. Hills & Day 1976). Some of the most massive stars known might in fact be merger remnants (Schneider et al. 2014) and also the Great Eruption of Car in the 1840s that created the Homunculus Nebula has been suggested to be the result of a merger event (e.g. Smith et al. 2018; Hirai et al. 2021). Some merged stars may also give rise to unusual supernova progenitors such as the surprising blue supergiant that lead to SN 1987A (Podsiadlowski & Joss 1989; Podsiadlowski et al. 1990, 1992; Morris & Podsiadlowski 2007; Menon & Heger 2017) while others are thought to be the progenitors of interacting and even superluminous supernovae (Justham et al. 2014). Still, the merging phase itself and also the evolution and final fates of merged stars leave many open questions. Schneider et al. (2025) present a general introduction to stellar mergers and the related common-envelope phase.
Mergers can be induced in various ways: by dynamical encounters in dense stellar systems such as star cluster (cf. “head-on collisions”), by orbital decay of binaries in dense interstellar/circumbinary media such as in AGN disks or during the formation of stars, by orbital dynamics in triple-star systems and other multiples (e.g. von Zeipel–Lidov–Kozai mechanism), and by mass transfer in isolated binary stars. The latter channel is supposed to lead to mergers for about 25% of all massive stars (; Sana et al. 2012) such that – when integrating over a constant star-formation rate in a stellar population like in the Milky Way – about 10% of all massive stars are expected to be remnants of mergers from isolated binary evolution (de Mink et al. 2014). Indeed, the binary evolution computations of Henneco et al. (2024) show that at least 12% of massive, mass-transferring binaries will merge, and this fraction goes up to if fully conservative mass transfer is assumed (Henneco 2025).
Ever since the discovery of surface magnetic fields in the A-type star 78 Vir by Babcock (1947), the origin of strong surface magnetism in massive OBA stars is debated. It has long been speculated that strong and large-scale magnetic fields are produced in stellar mergers and that they can persist over the long evolutionary timescales of stars (see e.g. Ferrario et al. 2009; Wickramasinghe et al. 2014). Today, direct simulations of essentialy every flavor of stellar mergers and other dynamic stellar phases show the efficient amplification of magnetic fields. This is the case in mergers of binary white dwarfs (e.g. Zhu et al. 2015; Pakmor et al. 2024), white dwarfs and neutron stars (e.g. Morán-Fraile et al. 2024), binary neutron stars (e.g. Kiuchi et al. 2024), low- and high-mass main-sequence (MS) stars (e.g. Schneider et al. 2019, hereafter Paper I; Ryu et al. 2024; Vynatheya et al. 2025) and also common-envelope events (e.g. Ohlmann et al. 2016b; Ondratschek et al. 2022; Vetter et al. 2024; Gagnier & Pejcha 2024).
The mergers of main-sequence (MS) stars offer a way to understand the origin of strong surface magnetic fields observationally found in about 10% of OBA stars (Landstreet 1992; Donati & Landstreet 2009; Fossati et al. 2015; Schöller et al. 2017; Grunhut et al. 2017). For example, a merger hypothesis has been put forward to explain the magnetic stars Sco (Schneider et al. 2016; Paper I), HR 2949 (Schneider et al. 2016), and HD 148937 (Frost et al. 2024). All three stars show clear rejuvenation that can be understood within a merger scenario, and HD 148937 is additionally surrounded by a bipolar nebula that appears to be the smoking gun for debris ejected in a merger event. In the case of Sco, we conduct three-dimensional (3D) magnetohydrodynamic (MHD) simulations of the merger of an initially and MS star and show strong magnetic field amplification and that the post-merger evolution of the star can explain many of the observed properties of Sco (Paper I). If the merger-generated magnetic fields can persist until the pre-supernova stage of Sco, the iron core-collapse of the progenitor likely leads to the formation of a highly magnetic neutron star, a so-called magnetar (Varma et al. 2023).
Here, we provide further analysis of the magnetic field amplification in the merger of the two MS stars originally presented in Paper I and the spin evolution of the merger remnant. In particular, we show how small-scale magnetic fields generated by the magneto-rotational and Kelvin–Helmholtz instabilities are converted into large-scale magnetic fields via large-scale fluid motions in the rotating post-merger star. We describe our computations in Sect. 2 and present our main findings in Sect. 3. The results are discussed in Sect. 4 and we conclude in Sect. 5.
2 Numerical methods
The initial conditions for the merger simulations in this paper are from stellar models computed with the stellar evolution code Mesa (Sect. 2.1) and represent a merger induced by binary star evolution. The hydrodynamics simulations of the merger are carried out using the moving-mesh magnetohydrodynamics code Arepo (Sect. 2.2).
2.1 Stellar evolution calculations
An and a MS star are evolved to an age of 9 Myr without rotation, using the stellar evolution code Mesa (version 9793, Paxton et al. 2011, 2013, 2015). The computations use an initial helium mass fraction of and solar-like metallicity (Asplund et al. 2009). We use the Ledoux criterion for convection and exponential convective core overshooting with a parameter of . Mixing-length theory (Henyey et al. 1965) is used with a mixing-length parameter of and a semi-convective efficiency of , but not applying the so-called MLT++ scheme in Mesa. We use the so-called “Dutch” wind mass-loss prescription with an efficiency factor of 1.0.
2.2 Hydrodynamics simulations
We carry out the hydrodynamics simulations using the finite-volume moving-mesh code Arepo (Springel 2010) with updates improving the convergence (Pakmor et al. 2016). We solve the ideal MHD equations using an HLLD Riemann solver (Miyoshi & Kusano 2005) and the Powell scheme for handling the divergence constraint (Powell et al. 1999); for more details and tests of the implementation, see Pakmor et al. (2011) and Pakmor & Springel (2013). We couple the OPAL equation of state (Rogers et al. 1996; Rogers & Nayfonov 2002) to the hydrodynamics which is the same as used in Mesa in the relevant part of the temperature–density plane.
For the merger simulations, we need to create initial conditions that consist of two stars that are about to merge. For these single stars, we generate stable models by following the method outlined in Ohlmann et al. (2017). We compute the stellar structure with Mesa and map it to a grid with cell centers on spherical shells distributed according to a HEALPix tesselation. Then, we run a relaxation because slight mismatches between pressure gradients and gravitional forces can induce spurious velocities that need to be damped away. We carry out the relaxation over ten dynamical timescales as given by the sound crossing time to ensure that the models remain stable. We give some parameters of the two initial stellar models in Table 1.
| Name | Mass | Radius | Luminosity | ||
| () | () | (d) | () | K | |
| MS-9M | 9.0 | 4.2 | 0.13 | 4.7 | 2.34 |
| MS-8M | 8.0 | 3.8 | 0.12 | 3.1 | 2.21 |
| a Sound crossing time | |||||
In a second step, the initial conditions for a merger are derived from these stable stellar models. In principle, they should be set up such that the stars just begin Roche lobe overflow. Such conditions, however, lead to a stable system that will not merge in the simulation for many orbits. In an additional step, initial conditions are generated that lead to a faster merger by adding an angular momentum loss term until a certain separation is reached. From there on, the simulation is continued without the loss term to capture the dynamics of the merger. The angular momentum loss term is implemented as an additional acceleration given as
where is a timescale on which the components merge and is the index of the cell. For this step, we take s and run the simulation for about 1.5 orbits with the loss term and continue without it from there on. For comparison, we start one model without the loss term after only half this time and it shows a very similar evolution. The simulations use 3.4 to 4 million cells in the high resolution runs and about 0.43 to 0.45 million cells in the low resolution runs. We set up the initial magnetic field as a dipole field, similar as in Ohlmann et al. (2016b), with a surface field of 1 G. We run one additional simulation without magnetic fields for comparison. We provide an overview of the models in Table 2.
| Model | Resolutiona | Initial separation | configuration |
| 1 | 3.4 – 4 | 6.4 | dipole |
| 2 | 0.43 | 6.4 | dipole |
| 3 | 0.45 | 7.0 | dipole |
| 4 | 0.43 | 6.4 | — |
| a Mean number of cells | |||
We can gain additional information on the merging process by tracking the origin of material during the merger. To this end, we add two scalars to the initial conditions that are passively advected with the flow. We set each of the scalars to 1 inside one of the stars and 0 everywhere else. Because we advect the scalars in the simulation, their final value gives the mass fraction of a cell originating from the corresponding star.
3 Results
A set of simulations has been run to study how the magnetic field is amplified in MS star mergers. These simulations use initial stellar models that are stable over a few dynamical timescales (Sect. 3.1). During the merger, the simulations show that the more massive primary star is disrupted around the secondary, forming a core–disk structure (Sect. 3.2). Magnetic fields are amplified in the simulations, first on small scales and later reaching large-scale structures (Sect. 3.3). These fields are mostly toroidal in the final merger product which shows a core–disk structure where the core originates mainly from the secondary and the disk mainly from the primary star (Sect. 3.4). During the evolution, the simulations show almost no ejection of mass (Sect. 3.5).
3.1 Initial models
Because the MS stars are quite compact compared to giants presented in Ohlmann et al. (2017), the method of Sec. 2.2 allows the generation of models that are stable in the hydrodynamical simulations for many dynamical timescales. This can be seen in Fig. 1, where spherically averaged profiles of the density, the Mach number, and the deviation from hydrostatic equilibrium are shown at different times during the relaxation run. One can see that the density profile (Fig. 1, top panel) remains the same in the interior of the star; only the boundary smears out slightly, which is to be expected due to the insufficient resolution in these low-density regions. For the same reason, the Mach numbers are slightly larger at the edge of the star, while the Mach number in the interior remains below (Fig. 1, middle panel). The hydrostatic equilibrium () is fulfilled to better than a few percent in most parts of the star (Fig. 1, bottom panel).
3.2 Evolution of the merger
The morphology of the flow during the merger is depicted in Fig. 3, showing slices of the density through the orbital plane and perpendicular to it, and in Fig. 4, showing slices of the passive scalar tracing material of the primary star through the orbital plane. Initially, mass transfer from the primary star to the secondary star sets in via the inner Lagrange point L1. This stream hits the secondary and leads to shear flows near the surface of the secondary. In addition, mass and also angular momentum are lost through the outer Lagrange points L2 and L3 (see, e.g., Fig. 3 at ), thereby shrinking the orbit. The orbital shrinkage accelerates until the stars merge: the initially more massive primary star is disrupted and distributed around the secondary star (see Figs. 3 and 4 at ). Directly after the merger, the cores of the stars are still separate (see central part of Fig. 4 at ). Shear in this central part causes hydrodynamical instabilities, mixing the cores and resulting in a symmetric merger product (Figs. 3 and 4 at ). Due to the large angular momentum of the binary, a disk-like structure forms around the rotating core with a mass of about (see Fig. 3 at in the – plane). The central core mainly originates from the less massive secondary, while the fraction of material from the primary increases outwards. The post-merger structure is described in more detail in Sect. 3.4.
The simulations of the merger start after the orbit of the binary system has shrunk in an additional step to (Model 1-3) or (Model 4), corresponding to orbital periods of about d and d, respectively. Strong mass transfer (about ) develops from the more massive primary star onto the secondary. The orbital separation decreases until the stars merge after roughly 23 (Model 1) to 35 (Model 4) orbits111For larger initial separations, the system becomes even more stable and takes more orbits to merge, rendering the simulations computationally unfeasible. (Fig. 2). As expected, Model 4 with a larger initial separation takes longer to merge than Models 1–3 with a smaller initial separation. Here, we define as the point in time where the orbital shrinkage becomes large, i.e., where the corresponding time scale becomes small and fulfills (where is the semi-major axis). Fig. 2 also shows that the distance varies with the orbital motions, i.e., the orbit is slightly eccentric. This is probably caused by the system not being perfectly Keplerian after the first step of artificially shrinking the system. The effect is smaller for the high-resolution Model 1.
3.3 Magnetic field amplification
During the rapid mass transfer leading up to the merger, magnetic fields are strongly amplified. Due to the complex flow patterns, it is difficult to unambiguously identify the amplification mechanisms. The maximum magnetic field strength grows exponentially for some time over several orders of magnitude until it saturates at about G, independent of the different initial conditions (Fig. 5, upper panel). Fits to an exponential function yield amplification times of 0.2 d to 1 d. The time evolution of the magnetic field in Fig. 6 (first row) shows that the initial amplification takes place in the region where the accretion stream hits the secondary. Strong shear flows induce Kelvin-Helmholtz instabilities, winding up the magnetic fields. The thickness of this layer is about .
The amplification continues in the mass stream from the primary to the secondary as it creates shear flows outside the Roche lobes of both stars (Fig. 6 at d). At this stage, we find that the flow is unstable to the magneto-rotational instability (MRI, Balbus & Hawley 1998). We estimate the length and time scale of the fastest growing MRI channel in our simulations (cf. Rembiasz et al. 2016): this yields a timescale of about 0.5 d and a wavelength of about which is similar to the scales in our simulation (time scale: 0.2–1 d, length scale:). Moreover, the strong amplification of the magnetic field starts as soon as the wavelength of the fastest growing MRI channel is resolved by a few hundred cells.
When the stellar cores merge, the magnetic field is distributed throughout the merger product, leading to a nearly symmetric distribution of the field strength (Fig. 6, d). Due to the differential rotation during the merger process, the field structure is mostly toroidal: the fraction of the toroidal magnetic energy is about of the total magnetic energy (for more details, see Sect. 3.4).
The amplification is slowing down and reaches saturation at later times, when the ratio of magnetic to turbulent kinetic energy approaches the percent level and comes closer to equipartition (Fig. 5, middle panel). As a proxy for the turbulent kinetic energy, we use the kinetic energy in the radial and direction, , which is not affected by the large-scale angular motion (see the comparison in Fig. 5, bottom panel).
The length scales can also be identified in power spectra of the kinetic and magnetic energy (Fig. 7). Most of the kinetic energy resides at large scales and low wave numbers for all times, as expected for the rotation of the stellar cores and of the post-merger product (Fig. 7, upper panel). The power spectrum follows a power law to smaller scales (exponent between and ), probably due to turbulence seeded by flow instabilities. The exponents are smaller in magnitude than the Kolmogorov exponent of that would be expected for a homogenous, isotropic turbulent cascade. Plots of power spectra of the magnetic energy (Fig. 7, lower panel), however, show that magnetic energy is injected at small scales (about –) and propagates to larger scales during the amplification (see d to d). After the cores have merged, redistribution of the magnetic energy leads to a similar spectrum at 5 d as for the kinetic energy, with most of the energy at large scales (about –). The exponents of the power law have been obtained by a fit to for between 0.1 and and are 1.38, , and for d, d, and d, respectively. This behavior is similar to the MRI simulations by Bhat et al. (2016), in which large-scale modes only grow late in the evolution when the amplification nears saturation. As mentioned there, large-scale magnetic fields seem to appear when the amplification process saturates (see also Ebrahimi et al. 2009).
The impact of the magnetic fields on the dynamics of the merger can be characterized by the ratio of magnetic pressure to gas pressure (see Fig. 6, bottom row). During the mass transfer leading up to the merger, stays below 1%, thus the effect of the magnetic fields on the dynamics of the merger is very small. Only after the merger, the ratio of the magnetic pressure rises in some parts up to values between 0.3 and 1, indicating that the magnetic fields have a stronger impact on the dynamics of the merger product, especially on the differential rotation.
3.4 Structure of the merger product
The large angular momentum in the binary system leads to the formation of a disk around a rotating core in the merger product (compare Fig. 3 at d). The density profile shows strong asymmetries between polar and equatorial regions indicating that the boundary between core and disk is located roughly at (Fig. 8). Moreover, the mass profile increases only in the equatorial region for radii above (Fig. 8, lower panel). Thus, almost all mass for radii above is contained in the equatorial region, i.e., in a disk-like structure. At d, the core contains a mass of about , and the disk a mass of about . Moreover, the disk comprises 58% of the total angular momentum.
Spherically averaged profiles of density, angular velocity, angular momentum, and magnetic field strength are shown in Fig. 9. The density profile remains very similar between d and d (Fig. 9, upper panel). At early times after the merger ( and d), the angular velocity profile reaches a maximum at some larger radius and is smaller in the core. Then, angular momentum redistribution leads to a nearly solid-body rotation in the central part () at d (Fig. 9, second row). In the outer part of the disk, however, no major angular momentum redistribution is visible (Fig. 9, third row)
The velocity at the core–disk boundary is about 700 km s-1 at d, i.e. 76% of the critical Keplerian rotational velocity. As can be seen in Fig. 9, second panel, the rotation curve in the disk always stays sub-Keplerian. After six days, this disk has completed about 170 rotations at the base and about 0.3 rotations at the edge. The ratio of kinetic energy to gravitational energy decreases from 0.07 at d to 0.05 at d and stays well below the limit of 0.14 for non-axisymmetric instabilities in rotating fluids (Ostriker & Tassoul 1969; Chandrasekhar 1970; Friedman & Schutz 1975; Lai & Shapiro 1995). Thus, no secular or dynamical, non-axisymmetric instabilities of the configuration are expected.
As already seen during the evolution of the merger (cf. Fig. 4), the very central part of the merger product originates from the secondary star, whereas the primary star is disrupted around the core. Only the inner of the product are dominated by material from the secondary, the rest of the core stems to similar parts from both stars. The disk, however, is predominantly formed from material of the primary star. This distribution is expected from the entropy and thus buoyancy profiles of the progenitor stars: due to the similar mass and hence also similar evolutionary stage, the entropy is similar and mixing is expected over large parts of the profile. Because the secondary has a lower mass, its entropy is slightly smaller and thus its material dominates the central part. This behaviour is similar to the case of direct collisions as explored by Glebbeek & Pols (2008) and Glebbeek et al. (2013), especially their case S.
The magnetic field structure of the final merger product is visualized in Fig. 10. The field structure is mostly toroidal; the energy of the toroidal magnetic field component is about 80-85% of the total magnetic energy. In this range, magnetic field configurations are expected to be stable in stars (Braithwaite & Spruit 2004; Braithwaite & Nordlund 2006a). Moreover, the configuration remains stable after the final merger and this time already corresponds to a few Alfvén timescales. However, in the regime of fast rotation (), the evolutionary timescale of the magnetic field is , where is the Alfvén timescale (Braithwaite & Cantiello 2013). Hence, great changes in the magnetic field are not expected to be seen over the timescales simulated here after the field has reached an equilibrium configuration. We expect Ohmic decay to operate on timescales that are comparable to the lifetime of the star (for more details, see Paper I). Hence, the magnetic field of the star should remain present at least until it reaches the main sequence. Modeling the subsequent stellar evolution in more detail shows that the merger product could evolve into a MS star with strong magnetic fields, similar to Sco (Paper I; Schneider et al. 2020).
The merger process leads to a denser and hotter core than a comparable 16.9 M⊙ MS star in full equilibrium: heat is mostly generated in the central region, leading to a larger internal energy in the core. The angular momentum redistribution leads to large rotational velocities outside the core (Fig. 9), thus the kinetic energy is larger compared to the internal energy outside the core, especially in the thick disk. This makes sense because the angular momentum of a binary increases outwards (), whereas the orbital energy decreases outwards (). This redistribution of energies is important to predict 1D merger products that can be used for further stellar evolution modeling (see also Sect. 4.2).
3.5 Dynamical ejecta
Because the further evolution of the merger product and also its optical display might be influenced by dynamical ejecta, we compute how much mass is unbound at the end of the simulation. To this end, we sum the mass of all cells in the simulation for which the total energy is positive. To get an upper limit, contributions from the potential, kinetic and also internal energies are taken into account. After the merger settled into the core–disk structure, at d, about of material are unbound, corresponding to 0.14% of the total mass of the system. Not much mass is ejected compared to other stellar mergers (e.g., common envelope phases, see Ohlmann et al. 2016a), also because the binding energy of the system is still quite large with about . The unbound mass moves outwards with an average radial velocity of about 180 km/s and is predominantly ejected near the orbital plane with a distribution that shows some tails to the polar directions.
4 Discussion
4.1 Impact of numerical parameters
The main numerical parameters that affect the simulation outcome are the resolution, the initial distance and the magnetic field configuration. To assess the impact of the resolution, we ran model 1 with about 10 times more cells than model 2 (cf. Table 2). Because the adaptive mesh leads to cells with nearly constant mass, this leads to a higher resolution in volume by a factor of 10, which roughly corresponds to a factor of 2 in each spatial dimension. Overall, these two simulations show quite similar behavior in their orbital evolution (Fig. 2). The evolution of the magnetic field amplification differs slightly for the two models (Fig. 6): the amplification sets in later for model 2 with less cells. The final field strength and configuration, however, are quite similar which demonstrates that they are robust with respect to the resolution.
The initial distance influences the outcome as more angular momentum is present in the system for larger initial distances. In principle, the simulation should start at Roche-lobe overflow; however, this is prohibitively expensive in computational time because the mass transfer progresses only slowly in the beginning, leading to a long time to the actual merger. As already described, we remove angular momentum artificially in order to obtain a starting configuration that leads to a merger more quickly. In reality, the system would have more angular momentum, leading probably to a longer mass loss via L2 and L3 in the phase prior to the actual merger. The largest impact of this is probably on the detailed structure of the outer part of the disk. To assess the impact of a larger initial separation, we ran model 3 with a larger initial distance. As one can see in Fig. 2, it takes longer until the cores finally merge, but the overall evolution is similar. Also, the magnetic field evolution (Fig. 6) shows a slightly different shape, but the final field strength and configuration are similar. Thus, we expect that the main conclusions are not affected by the detailed choice of this parameter, but that it mainly affects the outermost structure of the disk.
In addition, we ran a simulation without magnetic fields at all to assess their impact on the orbital evolution (model 4). The orbital evolution is very similar to model 2 with the same resolution (see Fig. 2), thus the impact of the magnetic fields on the dynamical evolution is small. Moreover, the ratio of magnetic pressure to gas pressure is only about 1% in the phase leading up to the merger. This indicates that the magnetic fields do not greatly influence the dynamics of the gas in this phase.
A typical numerical issue in MHD simulations is fulfilling the divergence constraint . In principle, the continuous equations of MHD preserve this constraint for initial conditions fulfilling it. However, discretization errors also lead to errors in . As explained in Pakmor & Springel (2013), we use the Powell method (Powell et al. 1999) to control the errors in the divergence. We have checked that the errors in our simulations behave very similarly as in the tests shown there (see Fig. 11): averages over larger regions are small and typically adjacent cells show different signs in the error, indicating that the large-scale impact of these local errors is small. The volume-weighted average of the relative divergence error is , , , and for the times given in Fig. 11.
4.2 Comparison to other simulations
Glebbeek et al. (2013) have conducted a series of stellar merger simulations as head-on collisions and subsequent stellar evolution calculations for a range of masses and evolutionary points. They used an SPH code with particles per star whereas we use a grid-based approach with 1 to 2 million cells per star. Moreover, we include MHD effects, which is the major novelty in our simulations. By simulating head-on collisions, they avoid dealing with rotation in the merger product and also with angular momentum transport. The collision may lead to more heating compared to the merger in our simulations. Moreover, there is probably no disk forming around the merger product in their simulations due to the missing angular momentum. In their simulations, the equal-mass collisions eject between 4% and 8% of the total mass (see their Figure 3), whereas in our simulation, only 0.14% of the mass is unbound (compare Sect. 3.5). This might be due to the difference in the initial conditions as material can escape more easily in head-on collisions, whereas in our merger simulations a large disk forms that is bound. In our simulation, the core is dominated by material from the secondary star, corresponding to “case S” in Glebbeek et al. (2013). As explained there, this can be attributed to the lower buoyancy in the secondary star. They find “case S” for relatively unevolved primaries at moderate mass ratios, which would also apply to our case.
Glebbeek & Pols (2008) have analyzed stellar collision products. However, they do not conduct hydrodynamic simulations, but they use the entropy sorting method from Lombardi et al. (2002) to compute the merger products and then they focus on the further evolution. Lombardi et al. (2002) use results from earlier collision simulations using an SPH code to construct a model for predicting the outcome of such a collision. This model mainly sorts the material by its entropy but takes into account more effects such as mass loss and shock heating. For our simulations, comparing the merger product with an entropy-sorted mix of the initial stellar models, one can clearly see similar effects: the entropy in the simulated merger is larger especially in the center and in the outermost parts. Simple entropy sorting of the initial stellar models does not provide the right structure of the merged stars. Shock heating, chemical mixing and changes in the total mass of the merged star and thereby in density and temperature modify the entropy of the original fluid elements. Still, the final merged star has an outward increasing entropy profile as is expected for buoyantly stable stratifications (Landau & Lifshitz 1987). The models from Lombardi et al. (2002) are based on earlier collision simulations, amongst others the off-axis collisions conducted by Sills et al. (2001). They use an SPH code with about particles for the full simulation and target low-mass stars, mainly to study blue stragglers similar to Lombardi et al. (2002) and Glebbeek & Pols (2008).
Ryu et al. (2024) simulate parabolic head-on collisions of low-mass stars with non-zero impact parameters similar to those mentioned above but they also use the moving-mesh MHD code Arepo and find strong magnetic field amplification. Such collisions contain less angular momentum than mergers in circular binaries studied here and thus form not so massive disk-like structures in the merger remnant as found here. The magnetic field amplification mechanism is otherwise in qualitative agreement with that found in our inspiral mergers of circular high-mass binaries: the collision mergers invoke several contact/grazing encounters before the final coalescence and in each phase small-scale dynamos amplify the magnetic fields. From the work presented in Ryu et al. (2024), however, it remains unclear whether also large-scale magnetic fields are formed. Vynatheya et al. (2025) compare head-on collisions of massive stars with binary mergers of the same stars. They find systematic differences between those, in particular they find that merger remnants have signs of an ordering of the magnetic field similar to our simulation, that is absent for collision remnants. They focus on a qualitative analysis of the final state of their collision and merger remnants. The latter seem consistent with our simulation.
It is remarkable that the magnetic field amplification process in mergers of MS stars is so similar to that in mergers of white dwarfs and neutron stars presented in Pakmor et al. (2024) and Kiuchi et al. (2024), respectively. Also in these binary white-dwarf and neutron-star mergers small scale magnetic fields are produced by a turbulent dynamo and magnetic energy is then transferred to larger scales by a large-scale, MRI-driven dynamo. A similar situation is also found in neutron star–white dwarf mergers (Morán-Fraile et al. 2024) and common-envelope phases of low-mass and high-mass stars (Ondratschek et al. 2022; Vetter et al. 2024). Despite the similarities, all these simulations feature outflows that are magnetically driven, whereas our simulations do not show such outflows. This might be because we could not follow our simulations for long enough after the merger.
5 Conclusions
We have studied in detail the magnetic-field amplification in the merger of a and main-sequence star in an isolated binary system and the properties of the resulting merged star. The simulations were first described in Paper I to explain observations of the magnetic massive star Sco. The focus of this study is on the magnetic field amplification process, the magnetic field geometry, the mixing during the merger and the rotational configuration of the resulting merger product. The main findings can be summarised as follows.
We see the formation of a star–torus structure in which the core of the initially less-massive secondary star is in the centre and the initially more-massive primary star is distributed around it. The material in the torus originates mostly from the primary. The rotationally-supported torus has a mass of and holds of the initial angular momentum of the binary.
There is hardly any dynamic mass loss and only of the total mass is ejected.
We observe robust magnetic-field amplification by more than ten orders of magnitude. The magnetic-field growth is initially sourced by turbulent energy in the merger and then by large-scale, ordered azimuthal fluid motions. In the end, the magnetic field energy reaches a super-equipartition level with respect to the kinetic energy in the radial and -direction, i.e. our proxy of the turbulent energy. The final outcome appears robust and does not depend on the initial separation of the binary, the numerical resolution and the strength of the seed magnetic field. We thus conclude that the memory of the initial seed magnetic field is lost.
The magnetic field is initially amplified by small-scale turbulent dynamos invoking the Kelvin-Helmholtz (shear) and magneto-rotational instabilities. Later, the large-scale, ordered azimuthal fluid motions drive another dynamo process that pumps the magnetic energy from small () to large () scales. The magnetic energy power spectrum in the final merger structure follows that of the kinetic energy and the merger remnant is threaded by a large-scale magnetic field.
The final magnetic-field geometry is characterised by a large-scale poloidal and toroidal structure with a still non-negligible small-scale component. The field structure resembles those found by Braithwaite & Spruit (2004) and Braithwaite & Nordlund (2006b) to be stable over thermal up to possibly nuclear timescales.
The spherically symmetric and pressure-supported central parts quickly approach solid-body rotation and transition into a Keplerian-like rotation profile in the torus.
This work is only the beginning of a detailed exploration of the complex physics in stellar mergers and many questions remain unanswered. The stability of the amplified magnetic fields needs to be assessed and the further evolution of the star–torus structure will determine the exact properties of merged stars such as their rotational velocities and surface chemical chemistry (see also Schneider et al. 2020). Addressing these questions is essential to further understand the properties of magnetic massive stars and their possible highly-magnetic white-dwarf and neutron-star remnants (cf. polars and magnetars).
Acknowledgements.
FRNS, FKR, RP, and VS were supported by the Klaus Tschira Foundation. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 945806) and is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1-390900948 (the Heidelberg STRUCTURES Excellence Cluster). The work of FKR is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – RO 3676/7-1, project number 537700965, and by the European Union (ERC, ExCEED, project number 101096243). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. For data processing and plotting, NumPy and SciPy (Oliphant 2007), IPython (Pérez & Granger 2007), and Matplotlib (Hunter 2007) were used.References
- Abbott et al. (2016) Abbott, B. P., Abbott, R., Abbott, T. D., et al. 2016, Phys. Rev. Lett., 116, 061102
- Abbott et al. (2017) Abbott, B. P., Abbott, R., Abbott, T. D., et al. 2017, Physical Review Letters, 119, 161101
- Asplund et al. (2009) Asplund, M., Grevesse, N., Sauval, A. J., & Scott, P. 2009, ARA&A, 47, 481
- Babcock (1947) Babcock, H. W. 1947, PASP, 59, 112
- Balbus & Hawley (1998) Balbus, S. A. & Hawley, J. F. 1998, Reviews of Modern Physics, 70, 1
- Bhat et al. (2016) Bhat, P., Ebrahimi, F., & Blackman, E. G. 2016, MNRAS, 462, 818
- Braithwaite & Cantiello (2013) Braithwaite, J. & Cantiello, M. 2013, MNRAS, 428, 2789
- Braithwaite & Nordlund (2006a) Braithwaite, J. & Nordlund, Å. 2006a, A&A, 450, 1077
- Braithwaite & Nordlund (2006b) Braithwaite, J. & Nordlund, Å. 2006b, A&A, 450, 1077
- Braithwaite & Spruit (2004) Braithwaite, J. & Spruit, H. C. 2004, Nature, 431, 819
- Braun & Langer (1995) Braun, H. & Langer, N. 1995, A&A, 297, 483
- Chandrasekhar (1970) Chandrasekhar, S. 1970, ApJ, 161, 561
- de Mink et al. (2014) de Mink, S. E., Sana, H., Langer, N., Izzard, R. G., & Schneider, F. R. N. 2014, ApJ, 782, 7
- Donati & Landstreet (2009) Donati, J.-F. & Landstreet, J. D. 2009, ARA&A, 47, 333
- Dray & Tout (2007) Dray, L. M. & Tout, C. A. 2007, MNRAS, 376, 61
- Ebrahimi et al. (2009) Ebrahimi, F., Prager, S. C., & Schnack, D. D. 2009, ApJ, 698, 233
- Ferrario et al. (2009) Ferrario, L., Pringle, J. E., Tout, C. A., & Wickramasinghe, D. T. 2009, MNRAS, 400, L71
- Fossati et al. (2015) Fossati, L., Castro, N., Schöller, M., et al. 2015, A&A, 582, A45
- Friedman & Schutz (1975) Friedman, J. L. & Schutz, B. F. 1975, ApJ, 199, L157
- Frost et al. (2024) Frost, A. J., Sana, H., Mahy, L., et al. 2024, Science, 384, 214
- Gagnier & Pejcha (2024) Gagnier, D. & Pejcha, O. 2024, A&A, 683, A4
- Glebbeek et al. (2013) Glebbeek, E., Gaburov, E., Portegies Zwart, S., & Pols, O. R. 2013, MNRAS, 434, 3497
- Glebbeek & Pols (2008) Glebbeek, E. & Pols, O. R. 2008, A&A, 488, 1017
- Grunhut et al. (2017) Grunhut, J. H., Wade, G. A., Neiner, C., et al. 2017, MNRAS, 465, 2432
- Hellings (1983) Hellings, P. 1983, Ap&SS, 96, 37
- Henneco (2025) Henneco, J. 2025, PhD thesis, Heidelberg University
- Henneco et al. (2024) Henneco, J., Schneider, F. R. N., & Laplace, E. 2024, A&A, 682, A169
- Henyey et al. (1965) Henyey, L., Vardya, M. S., & Bodenheimer, P. 1965, ApJ, 142, 841
- Hills & Day (1976) Hills, J. G. & Day, C. A. 1976, Astrophys. Lett., 17, 87
- Hirai et al. (2021) Hirai, R., Podsiadlowski, P., Owocki, S. P., Schneider, F. R. N., & Smith, N. 2021, MNRAS, 503, 4276
- Hunter (2007) Hunter, J. D. 2007, Computing in Science & Engineering, 9, 90
- Justham et al. (2014) Justham, S., Podsiadlowski, P., & Vink, J. S. 2014, ApJ, 796, 121
- Kiuchi et al. (2024) Kiuchi, K., Reboul-Salze, A., Shibata, M., & Sekiguchi, Y. 2024, Nature Astronomy, 8, 298
- Lai & Shapiro (1995) Lai, D. & Shapiro, S. L. 1995, ApJ, 442, 259
- Landau & Lifshitz (1987) Landau, L. D. & Lifshitz, E. M. 1987, Fluid Mechanics (Course of Theoretical Physics: Volume 6), 2nd edn. (Oxford: Butterworth-Heinemann)
- Landstreet (1992) Landstreet, J. D. 1992, A&A Rev., 4, 35
- Lombardi et al. (2002) Lombardi, Jr., J. C., Warren, J. S., Rasio, F. A., Sills, A., & Warren, A. R. 2002, ApJ, 568, 939
- Menon & Heger (2017) Menon, A. & Heger, A. 2017, MNRAS, 469, 4649
- Miyoshi & Kusano (2005) Miyoshi, T. & Kusano, K. 2005, Journal of Computational Physics, 208, 315
- Morán-Fraile et al. (2024) Morán-Fraile, J., Röpke, F. K., Pakmor, R., et al. 2024, A&A, 681, A41
- Morris & Podsiadlowski (2007) Morris, T. & Podsiadlowski, P. 2007, Science, 315, 1103
- Munari et al. (2002) Munari, U., Henden, A., Kiyota, S., et al. 2002, A&A, 389, L51
- Ohlmann et al. (2016a) Ohlmann, S. T., Röpke, F. K., Pakmor, R., & Springel, V. 2016a, ApJ, 816, L9
- Ohlmann et al. (2017) Ohlmann, S. T., Röpke, F. K., Pakmor, R., & Springel, V. 2017, A&A, 599, A5
- Ohlmann et al. (2016b) Ohlmann, S. T., Röpke, F. K., Pakmor, R., Springel, V., & Müller, E. 2016b, MNRAS, 462, L121
- Oliphant (2007) Oliphant, T. E. 2007, Computing in Science & Engineering, 9, 10
- Ondratschek et al. (2022) Ondratschek, P. A., Röpke, F. K., Schneider, F. R. N., et al. 2022, A&A, 660, L8
- Ostriker & Tassoul (1969) Ostriker, J. P. & Tassoul, J. L. 1969, ApJ, 155, 987
- Pakmor et al. (2011) Pakmor, R., Bauer, A., & Springel, V. 2011, MNRAS, 418, 1392
- Pakmor et al. (2024) Pakmor, R., Pelisoli, I., Justham, S., et al. 2024, A&A, 691, A179
- Pakmor & Springel (2013) Pakmor, R. & Springel, V. 2013, MNRAS, 432, 176
- Pakmor et al. (2016) Pakmor, R., Springel, V., Bauer, A., et al. 2016, MNRAS, 455, 1134
- Paxton et al. (2011) Paxton, B., Bildsten, L., Dotter, A., et al. 2011, ApJS, 192, 3
- Paxton et al. (2013) Paxton, B., Cantiello, M., Arras, P., et al. 2013, ApJS, 208, 4
- Paxton et al. (2015) Paxton, B., Marchant, P., Schwab, J., et al. 2015, ApJS, 220, 15
- Pérez & Granger (2007) Pérez, F. & Granger, B. E. 2007, Computing in Science & Engineering, 9, 21
- Podsiadlowski & Joss (1989) Podsiadlowski, P. & Joss, P. C. 1989, Nature, 338, 401
- Podsiadlowski et al. (1992) Podsiadlowski, P., Joss, P. C., & Hsu, J. J. L. 1992, ApJ, 391, 246
- Podsiadlowski et al. (1990) Podsiadlowski, P., Joss, P. C., & Rappaport, S. 1990, A&A, 227, L9
- Powell et al. (1999) Powell, K. G., Roe, P. L., Linde, T. J., Gombosi, T. I., & de Zeeuw, D. L. 1999, Journal of Computational Physics, 154, 284
- Rembiasz et al. (2016) Rembiasz, T., Obergaulinger, M., Cerdá-Durán, P., Müller, E., & Aloy, M. A. 2016, MNRAS, 456, 3782
- Rogers & Nayfonov (2002) Rogers, F. J. & Nayfonov, A. 2002, ApJ, 576, 1064
- Rogers et al. (1996) Rogers, F. J., Swenson, F. J., & Iglesias, C. A. 1996, ApJ, 456, 902
- Ryu et al. (2024) Ryu, T., Sills, A., Pakmor, R., de Mink, S., & Mathieu, R. 2024, arXiv e-prints, arXiv:2410.00148
- Sana et al. (2012) Sana, H., de Mink, S. E., de Koter, A., et al. 2012, Science, 337, 444
- Schneider et al. (2014) Schneider, F. R. N., Izzard, R. G., de Mink, S. E., et al. 2014, ApJ, 780, 117
- Schneider et al. (2025) Schneider, F. R. N., Lau, M. Y. M., & Roepke, F. K. 2025, arXiv e-prints, arXiv:2502.00111
- Schneider et al. (2020) Schneider, F. R. N., Ohlmann, S. T., Podsiadlowski, P., et al. 2020, MNRAS, 495, 2796
- Schneider et al. (2019) Schneider, F. R. N., Ohlmann, S. T., Podsiadlowski, P., et al. 2019, Nature, 574, 211
- Schneider et al. (2016) Schneider, F. R. N., Podsiadlowski, P., Langer, N., Castro, N., & Fossati, L. 2016, MNRAS, 457, 2355
- Schöller et al. (2017) Schöller, M., Hubrig, S., Fossati, L., et al. 2017, A&A, 599, A66
- Sills et al. (2001) Sills, A., Faber, J. A., Lombardi, Jr., J. C., Rasio, F. A., & Warren, A. R. 2001, ApJ, 548, 323
- Smith et al. (2018) Smith, N., Andrews, J. E., Rest, A., et al. 2018, MNRAS, 480, 1466
- Springel (2010) Springel, V. 2010, MNRAS, 401, 791
- Tylenda et al. (2011) Tylenda, R., Hajduk, M., Kamiński, T., et al. 2011, A&A, 528, A114
- Varma et al. (2023) Varma, V., Müller, B., & Schneider, F. R. N. 2023, MNRAS, 518, 3622
- Vetter et al. (2024) Vetter, M., Röpke, F. K., Schneider, F. R. N., et al. 2024, A&A, 691, A244
- Vynatheya et al. (2025) Vynatheya, P., Ryu, T., Wang, C., Sills, A., & Pakmor, R. 2025, arXiv e-prints, arXiv:2510.13736
- Wickramasinghe et al. (2014) Wickramasinghe, D. T., Tout, C. A., & Ferrario, L. 2014, MNRAS, 437, 675
- Zhu et al. (2015) Zhu, C., Pakmor, R., van Kerkwijk, M. H., & Chang, P. 2015, ApJ, 806, L1